Abbeville Fixture Company manufactures units in a small manufacturing facility. The units are made from brass. Manufacturing has 60 employees. Each employee presently provides 36 hours of labor per week. Information about a production week is as follows:

Standard wage per hour $17.4
Standard labor time per unit 20 min.
Standard number of lbs. of brass 1.9 lbs.
Standard price per lb. of brass $9.25
Actual price per lb. of brass $9.5
Actual lbs. of brass used during the week 17,613 lbs.
Number of units produced during the week 9,000
Actual wage per hour $17.92
Actual hours for the week (60 employees × 36 hours) 2,160 hrs.

Required:

a.  Determine the standard cost per unit for direct materials and direct labor. Round the cost per unit to two decimal places.

Direct materials standard cost per unit $
Direct labor standard cost per unit $
Total standard cost per unit $

b.  Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Materials Price Variance $
Direct Materials Quantity Variance $
Total Direct Materials Cost Variance $

c.  Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance $
Direct Labor Time Variance $
Total Direct Labor Cost Variance $
Answer #1
a
Direct materials standard cost per unit 17.58 =1.9*9.25
Direct labor standard cost per unit 5.80 =17.4*20/60
Total standard cost per unit 23.38
b
Direct Materials Price Variance 4403 Unfavorable =17613*(9.5-9.25)
Direct Materials Quantity Variance 4745 Unfavorable =9.25*(17613-9000*1.9)
Total Direct Materials Cost Variance 9149 or 9148 Unfavorable =(17613*9.5)-(9000*1.9*9.25)
c
Direct Labor Rate Variance 1123 Unfavorable =2160*(17.92-17.4)
Direct Labor Time Variance -14616 Favorable =17.4*(2160-9000*20/60)
Total Direct Labor Cost Variance -13493 Favorable =(2160*17.92)-(9000*17.4*20/60)